OpenStudy (prince1342):

Given a soda can with a volume of 15 and a diameter of 2, what is the volume of a cone that fits perfectly inside the soda can? (Hint: only enter numerals in the answer blank).

OpenStudy (prince1342):

@563blackghost Help please

OpenStudy (prince1342):

The formula of a cylinder is pi times heght and radius but I cant seem to figure the height

563blackghost (563blackghost):

this one is kinda tricky for me so one sec .....

OpenStudy (prince1342):

Alright ^_^

rebeccaxhawaii (rebeccaxhawaii):

so a cylinder and cone

563blackghost (563blackghost):

Ok we first must find the height of the cylinder so we input what we know.... \(\huge{15=(3.14)1^{2}(h)}\) So we multiply \(\large{3.14}\) by \(\large{1^{2}}\)... \(\huge\color{orange}{15=3.14(h)}\) in which we divide \(\large{15}\) by \(\large{3.14}\).... \(\huge\color{purple}{h=4.777}\) Now that we know of this we can find the cones volume...

OpenStudy (prince1342):

Ohhh okay, I'm with you so far

OpenStudy (mathmale):

Suggestions: 1. Sketch the cylinder and cone separately. Label everything. 2. Draw the cylinder again, this time with the cone inside. 3. Determine how the height of the cone relates to the height of the cylinder. 4. Determine how the area of the base of the cylinder relates to that of the cone. 5. Determine the radius of the cylinder from the given info. 6. Determine the radius of the cone's base.

OpenStudy (prince1342):

I got 14.99978 for the cone

OpenStudy (mathmale):

7. Write out both the formula for the volume of a cone and that for the volume of a cylinder. 8. Compare the two formulas. Anything useful there?

OpenStudy (mathmale):

@Prince1342: Please share all of y our work. It's hard to give you feedback when you provide nothing but an "answer." But I can tell you that if the vol. of the cylinder is 15, the vol. of the cone is a lot less.

563blackghost (563blackghost):

So the formula for volume of a cone is \(\huge{V=\frac{1}{3} \pi r^{2}h}\) in which we input the radius which is half of \(\large{2}\) so it would be \(\large{1}\) and the height would be \(\large{4.77}\)... \(\huge{V=\frac{1}{3}(3.14)(1^{2})4.77}\) So we now simplify...

OpenStudy (prince1342):

@mathmale I know I'm about to do that actually haha

OpenStudy (mathmale):

@Prince1342 : In your shoes I'd summarize all of the info shared with you by other users and then pick and choose from it: Find the most relevant and most likely correct info and work with that.

OpenStudy (prince1342):

When plugging in those equations I got 5 (4.99999) and @mathmale that's a good point, thanks!

563blackghost (563blackghost):

That would be correct @Prince1342 ^^

OpenStudy (prince1342):

Wow awesome! Thank you both, I really appreciate it

OpenStudy (mathmale):

Good luck. Unfortunately, I need to get off the 'Net now. Hope you'll post sketchs of the cylinder - cone combo and label everything. Note that the cyl. and the cone have the same base area and the same height. Generally, a cone has volume V = (1/3)*base*height. This is a great shortcut to the answer you need.

OpenStudy (mathmale):

So your 5 is correct. But...that's 5 what? Units of measurement are important.

OpenStudy (prince1342):

5 cm

OpenStudy (prince1342):

That is indeed a good short cut!!1

563blackghost (563blackghost):

@mathmale is right units of measure are important even if it is as a mutliple choice they can get you with either squared or cubed...

OpenStudy (prince1342):

There is one more question I need help on, I'll post it seperately and tag you is that okay?

563blackghost (563blackghost):

Sure ^^